A particular fruit's weights are normally distributed, with a mean of 657 grams and a standard deviation of 38 grams. The heaviest 15% of fruits weigh more than how many grams? Give your answer to the nearest gram.
Answer #1

Solution:-

Given that,

mean = = 657 grams

standard deviation = = 38 grams

Using standard normal table,

P(Z > z) = 15%

= 1 - P(Z < z) = 0.15  

= P(Z < z) = 1 - 0.15

= P(Z < z ) = 0.85

= P(Z < 1.036 ) = 0.85  

z = 1.036

Using z-score formula,

x = z * +

x = 1.036 * 38 + 657

x = 696.37

15% haviest weigh is 696 grams